# Parameter-Efficient Fine-Tuning of DINOv2 for Large-Scale Font Classification

### v1

@@ Abstract
We present a font classification system capable of identifying 394 font families from rendered text images. Our approach fine-tunes a DINOv2 Vision Transformer using Low-Rank Adaptation (LoRA), achieving approximately 86% top-1 accuracy while training fewer than 1% of the model’s 87.2M parameters. We introduce a synthetic dataset generation pipeline that renders Google Fonts at scale with diverse augmentations including randomized colors, alignment, line wrapping, and Gaussian noise, producing training images that generalize to real-world typographic samples. The model incorporates built-in preprocessing to ensure consistency between training and inference, and is deployed as a HuggingFace Inference Endpoint. We release the model, dataset, and full training pipeline as open-source resources.
@@ 1 Introduction
In this work, we take advantage of recent advances in self-supervised visual representation learning to address these challenges. Specifically, we fine-tune DINOv2 (Oquab et al., 2024), a Vision Transformer pre-trained with self-distillation on a large curated image dataset, using Low-Rank Adaptation (LoRA) (Hu et al., 2022). This parameter-efficient strategy allows us to adapt a powerful general-purpose vision backbone to the font classification domain while training only a small fraction of the total parameters.
@@
    A parameter-efficient fine-tuning approach for font classification that adapts DINOv2 with LoRA, achieving 86% top-1 accuracy across 394 font classes.
@@
    A scalable synthetic dataset generation pipeline that renders Google Fonts with diverse augmentations to produce training data robust to visual variation.
@@
    An end-to-end system with built-in preprocessing that ensures consistency between training and deployment, released as an open-source model and dataset on HuggingFace.
This model serves as a component of Anything, [^1] https://anything.com which achieved the #1 ranking on Design Arena’s builders leaderboard, [^2] https://www.designarena.ai/leaderboard/builder a crowdsourced benchmark for AI-generated design with over 150,000 voters.
@@ 3.1 Font Selection
We source fonts from the Google Fonts open-source repository. [^3] https://github.com/google/fonts From the full catalog, we select a curated set of 31 base font families spanning diverse typographic categories: sans-serif (e.g., Inter, Roboto, Poppins, DM Sans), serif (e.g., Crimson Pro, Playfair Display, Merriweather), monospace (JetBrains Mono), and display faces (e.g., Ultra, Big Shoulders Text). Selection was guided by three criteria: (1) prevalence in modern design tooling—we prioritized fonts frequently used in Figma, Dribbble, and similar platforms for website and mobile design; (2) visual distinctiveness—we excluded near-duplicate families that differ only in minor metrics to keep classes well-separated; and (3) representativeness of high-quality digital design—the set reflects the typographic choices typical of professional web and mobile interfaces. Many of these families include variable font axes (primarily weight), and we enumerate each axis variation as a separate class. This produces a total of 394 distinct font variants for classification.
@@ 3.1.1 Iterative Font Curation
The final font set is the result of four iterations of the model and dataset, each informed by the failures of the previous version:
  • v1 (10 families, 102 variants, 91.6% accuracy). [^4] https://huggingface.co/dchen0/font-classifier A small proof-of-concept trained on a handful of families (Inter, Roboto, Poppins, and several others). High accuracy, but the narrow label space did not cover enough fonts to be useful in practice.
  • v2 (many dozens of families, 700 variants). [^5] https://huggingface.co/dchen0/font-classifier-v2 A broad expansion that included a large portion of the Google Fonts catalog. The much larger label space introduced many visually near-identical font pairs (e.g., families that differ only in minor spacing or hinting), causing widespread confusion between classes and degrading accuracy.
  • v3 (20 families, 300 variants, 59% accuracy). [^6] https://huggingface.co/dchen0/font-classifier-v3 Scaled back from v2 by removing duplicate-like families, but the training dataset was too small (approximately 9,400 images) relative to the label count, and several included families still overlapped visually. Accuracy dropped significantly.
  • v4 (31 families, 394 variants, 86% accuracy). [^7] https://huggingface.co/dchen0/font_classifier_v4 The current version. The font set was rebuilt from scratch using the three selection criteria described above. The training dataset was scaled up substantially (approximately 575 images per variant). A critical engineering fix—embedding preprocessing into the model’s forward method—eliminated train-serve skew that had silently hurt earlier versions.
@@ 3.2 Synthetic Image Generation
    Rendering. Text is rendered at 1024px font size onto a canvas with padding on all sides (default 50px) using the Pillow library. The canvas is then cropped to the tight bounding box of the rendered glyphs (i.e., all empty space beyond the ink is removed). The cropped glyph image is resized so that its height equals 224 pixels, with the width scaled proportionally to preserve the aspect ratio, using Lanczos resampling.
@@
The resulting dataset, font_crops_v4, [^10] https://huggingface.co/datasets/dchen0/font_crops_v4 contains approximately 575 training images per font variant, with a held-out test set of 40 images per variant.
@@ 4.1 Model Architecture
Our model builds on the DINOv2-Base architecture (Oquab et al., 2024), specifically the dinov2baseimagenet1k1layer variant pre-trained on ImageNet-1K. This model uses the standard ViT-B/14 architecture with patch size 14, embedding dimension 768, 12 attention heads, and 12 transformer blocks.
@@ 4.2 LoRA Configuration
This configuration adds approximately 150K trainable parameters on top of the 87M frozen backbone parameters, representing less than 0.2% of the total model size.
@@ 4.3 Preprocessing
To eliminate preprocessing discrepancies, we implement a custom FontClassifierWithPreprocessing class that overrides the model’s forward method to apply preprocessing within the model itself. This ensures that any deployment path—HuggingFace pipeline, inference endpoint, or direct model call—applies identical transformations.
@@ 5.2 Infrastructure
Training is conducted using PyTorch with HuggingFace Transformers and the PEFT library for LoRA integration. The training script automatically selects the best available hardware accelerator: Apple MPS (Metal Performance Shaders, for Apple Silicon GPUs) is preferred if available, followed by CUDA (for NVIDIA GPUs), with CPU as a fallback. Mixed-precision training (FP16) is enabled on CUDA to reduce memory usage and training time; MPS and CPU use full FP32 precision. TensorBoard is used for monitoring training and validation metrics in real time.
@@ 6.1 Classification Accuracy
The fine-tuned model achieves approximately 86% top-1 accuracy on the held-out test set across 394 font classes. Given the fine-grained nature of the task—many classes differ only in weight (e.g., Roboto-300 vs. Roboto-400) or subtle stylistic features—this represents strong performance.
@@ 6.2 Analysis
  • Weight variant confusion. The most common errors occur between adjacent weight variants of the same font family (e.g., weight 400 vs. 500). This is expected, as the visual difference between adjacent weights can be as small as a few pixels of stroke width at the $224\times 224$ resolution.
@@
Figure 2: Row-normalized confusion matrix across all font classes, grouped by font family. The strong diagonal indicates high per-class accuracy, with most off-diagonal mass concentrated within family blocks (weight variant confusion).
Figure 3: Top-20 most frequent misclassification pairs. The majority involve adjacent weight variants within the same font family, confirming that inter-family classification is near-perfect.
@@
Figure 5: Classification accuracy broken down by font family, sorted from lowest to highest. Most families achieve near-perfect accuracy; lower-performing families tend to have many visually similar weight variants.
Figure 6: Per-family accuracy plotted against the number of weight variants in each family. Families with more variants (and thus more visually similar classes to distinguish) tend to have lower accuracy, consistent with the hypothesis that intra-family weight confusion drives most classification errors.
Figure 7: Hierarchical clustering (UPGMA) of 31 font families by cosine distance between mean [CLS] embeddings. The model’s learned representations recover typographic structure: serif, sans-serif, monospace, and display families occupy distinct branches.
@@
For a continuous measure, we define a severity-weighted error rate (SWER) using the model’s own learned representations. For each of the $K=394$ classes, we compute a centroid $\mathbf{c}_{k}$ by averaging the [CLS] token embeddings (from the final hidden layer) of all test samples belonging to class $k$. The pairwise cosine distance between centroids is:
$d(i,j)=1-\frac{\mathbf{c}_{i}\cdot\mathbf{c}_{j}}{\|\mathbf{c}_{i}\|\,\|\mathbf{c}_{j}\|}$ (2)
@@
Table 2: Error severity analysis. Family-level accuracy collapses weight variants; relative severity $\Pi$ is a scale-invariant ratio of the model’s embedding-distance-weighted error to the random baseline.
@@
Top-1 accuracy (variant-level) 38.5%
Family-level accuracy 40.2%
SWER (model) 0.3877
SWER (random baseline) 0.7008
Relative severity $\Pi$ 0.5532
Figure 7 visualizes the font family similarity hierarchy recovered by UPGMA clustering on the centroid distances. Serif, sans-serif, monospace, and display families occupy distinct branches, confirming that the embedding space captures typographic structure.
@@ 7 Deployment
The trained model is released on HuggingFace Hub as dchen0/font_classifier_v4.111111https://huggingface.co/dchen0/font_classifier_v4 Before upload, the LoRA adapter weights are merged into the base model using PEFT’s merge_and_unload() method. Concretely, each low-rank update $\frac{\alpha}{r}BA$ is added directly to the corresponding frozen weight matrix $W$, producing a single set of full-rank weights that behave identically to the adapted model but no longer require the PEFT library at inference time.
The model supports three deployment modes:
  1. 1.
    HuggingFace Inference Endpoints. A custom handler.py accepts raw bytes, base64-encoded strings, or PIL images, applies standardized preprocessing, and returns top-5 predictions with softmax confidence scores.
  2. 2.
    HuggingFace Pipeline. Standard pipeline("image-classification") usage with automatic model and processor loading.
  3. 3.
    Local inference. A command-line tool (serve_model.py) for single-image prediction with top-5 output.
@@ 8 Discussion
Natural extensions include: (1) training on a broader set of fonts beyond the current 31 families, (2) incorporating real-world training images via domain adaptation, (3) jointly predicting font family and attributes (weight, width, style) in a multi-task framework, and (4) integrating text detection to enable end-to-end font recognition from uncropped images.
@@ 9 Conclusion
We have presented a font classification system that identifies 394 font variants with 86% accuracy by fine-tuning a DINOv2 Vision Transformer with LoRA. The approach is parameter-efficient, training fewer than 0.2% of model parameters, and is supported by a reproducible synthetic data generation pipeline. The model, dataset, and code are publicly available to support further research in document analysis and typographic recognition. [^12] Source code: https://github.com/Create-Inc/font-model

### v2

@@ Abstract
[…]
@@ 1 Introduction
Existing font recognition benchmarks such as AdobeVFR (Wang et al., 2015) focus almost exclusively on commercial typefaces. No public benchmark covers the open-source fonts that dominate modern web and mobile design—Google Fonts alone powers over 70 billion monthly font views. This gap leaves practitioners without a standardized way to evaluate font classification systems for the fonts they actually encounter.
In this work, we address both the benchmark gap and the classification task. We introduce GoogleFontsBench, a public benchmark for open-source font classification comprising 394 font variants across 32 families, a reproducible synthetic data generation pipeline, and a typographically-grounded evaluation protocol. We establish baselines on this benchmark using six fine-tuning strategies, finding that parameter-efficient adaptation of DINOv2 (Oquab et al., 2024) with LoRA (Hu et al., 2022) achieves the strongest efficiency–accuracy trade-off.
@@
    GoogleFontsBench: the first public font classification benchmark for open-source web fonts, with 394 classes, $\sim$}225K training images, and a standardized evaluation protocol including a typographic-distance-based severity metric (SWER).
@@
    Baseline results from six fine-tuning strategies (linear probe, LoRA at three ranks, full fine-tuning, and ResNet-50), establishing reference points for future work.
@@
[…]
  4. 4.
    An end-to-end system with built-in preprocessing that ensures consistency between training and deployment, released as open-source on HuggingFace.
This model is deployed as a component of a production design tool. [^1] https://anything.com
@@ 3.1 Font Selection
We source fonts from the Google Fonts open-source repository. [^2] https://github.com/google/fonts From the full catalog, we select a curated set of 32 base font families spanning diverse typographic categories: sans-serif (e.g., Inter, Roboto, Poppins, DM Sans), serif (e.g., Crimson Pro, Playfair Display, Merriweather), monospace (JetBrains Mono), and display faces (e.g., Ultra, Big Shoulders Text). Selection was guided by three criteria: (1) prevalence in modern design tooling—we selected fonts commonly encountered in Figma, Dribbble, and similar platforms, based on the authors’ experience building production design tools for website and mobile design; (2) visual distinctiveness—we excluded near-duplicate families that differ only in minor metrics to keep classes well-separated; and (3) representativeness of high-quality digital design—the set reflects the typographic choices typical of professional web and mobile interfaces. Many of these families include variable font axes (primarily weight), and we enumerate each axis variation as a separate class. This produces a total of 394 distinct font variants for classification.
@@ 3.1.1 Iterative Font Curation
The final font set is the result of four dataset iterations, each informed by the failures of the previous version. An initial proof-of-concept with 10 families (102 variants) achieved high accuracy but covered too few fonts to be practical. Expanding aggressively to over 700 variants from the broader Google Fonts catalog introduced many visually near-identical pairs (families differing only in minor spacing or hinting), which degraded accuracy substantially. Scaling back to 20 families helped but exposed insufficient training data and residual visual overlaps. The current version rebuilds the font set from scratch using the selection criteria described above, settling on 32 families with 394 variants and approximately 575 training images per variant. A critical engineering fix in this iteration—embedding preprocessing into the model’s forward method—eliminated train–serve skew that had silently hurt earlier versions. All dataset versions are publicly available on HuggingFace.
@@ 3.2 Synthetic Image Generation
    Rendering. Text is rendered at 1024px font size onto a canvas with padding on all sides (128px) using the Pillow library. The canvas is then cropped to the tight bounding box of the rendered glyphs (i.e., all empty space beyond the ink is removed). The cropped glyph image is resized so that its height equals 224 pixels, with the width scaled proportionally to preserve the aspect ratio, using Lanczos resampling.
@@
The resulting dataset, font_crops_v5, [^3] https://huggingface.co/datasets/dchen0/font_crops_v5 contains approximately 575 training images per font variant, with a held-out test set of 40 images per variant.
@@ 4.1 Model Architecture
Our model builds on the DINOv2-Base architecture (Oquab et al., 2024), specifically the dinov2-base-imagenet1k-1-layer variant pre-trained on ImageNet-1K. This model uses the standard ViT-B/14 architecture with patch size 14, embedding dimension 768, 12 attention heads, and 12 transformer blocks.
@@ 4.2 LoRA Configuration
This configuration adds approximately 295K trainable LoRA parameters. Together with the 606K-parameter classification head (which is also trainable), the total number of trainable parameters is approximately 900K, or roughly 1% of the 87M frozen backbone.
@@ 4.3 Preprocessing
To eliminate preprocessing discrepancies, we define a shared transform pipeline (get_inference_transform in handler.py) that is applied before every model call. The EndpointHandler class used for HuggingFace Inference Endpoints applies this same pipeline, ensuring that any deployment path—inference endpoint, batch evaluation, or direct model call—uses identical transformations.
@@ 5.2 Infrastructure
Training is conducted using PyTorch with HuggingFace Transformers and the PEFT library for LoRA integration on Vast.ai cloud GPU instances equipped with NVIDIA RTX 3090 GPUs (24 GB VRAM). Mixed-precision training (FP16) is enabled to reduce memory usage and training time. With the LoRA r=8r{=}8 configuration and a batch size of 64, each epoch takes approximately 20 minutes; total training time for 100 epochs is approximately 33 hours per configuration.
@@ 6.1 Classification Accuracy
The best-performing configuration (LoRA r=8r{=}8) achieves 99.0% top-1 accuracy on the held-out test set across 394 font classes (Table 2). LoRA r=8r{=}8 achieved a peak validation accuracy of 99.0% at epoch 99 after full 100-epoch training, marginally outperforming r=16r{=}16 which peaked at 98.9%. At a fixed 50-epoch budget (Table 3), the two configurations perform identically at 98.7%. Wilson score 95% confidence intervals (Table 3) confirm that the differences between LoRA ranks are not statistically significant, suggesting rank has minimal impact beyond r=4r{=}4 for this task. Given the fine-grained nature of the task—many classes differ only in weight (e.g., Roboto-300 vs. Roboto-400) or subtle stylistic features—this represents strong performance.
@@ 6.2 Analysis
  • Weight variant confusion. The most frequent individual error pairs remain within-family weight confusions (e.g., weight 400 vs. 500), though at 99% accuracy the error distribution also includes cross-family pairs with similar visual characteristics. This is expected, as the visual difference between adjacent weights can be as small as a few pixels of stroke width at the $224\times 224$ resolution.
@@
Figure 2: Row-normalized confusion matrix across all font classes, grouped by font family. The strong diagonal indicates high per-class accuracy, with remaining off-diagonal mass primarily within family blocks (weight variant confusion).
Figure 3: Top-20 most frequent misclassification pairs. At 99% top-1 accuracy, the remaining misclassifications are diverse—including both within-family weight confusions and cross-family pairs with similar visual characteristics (e.g., thin sans-serif variants).
@@
Figure 5: Top-1 variant-level accuracy broken down by font family, sorted from lowest to highest. Most families achieve near-perfect accuracy; lower-performing families tend to have many visually similar weight variants.
Figure 6: Top-1 variant-level accuracy plotted against the number of weight variants in each family. At 99% overall accuracy, per-family accuracy is uniformly high (0.96–1.00), though a slight downward trend remains for families with many visually similar weight variants. Dashed line shows an OLS linear fit.
Figure 7: Hierarchical clustering (UPGMA) of 32 font families by cosine distance between mean [CLS] embeddings. The model’s learned representations recover typographic structure: serif, sans-serif, monospace, and display families occupy distinct branches.
@@
For a continuous measure, we define a severity-weighted error rate (SWER) using a typographic metadata distance that is independent of the model being evaluated. The pairwise distance $d(i,j)$ between font classes $i$ and $j$ is defined by a tier structure reflecting typographic similarity:
  • $d=0$: same variant (correct prediction).
  • $d=0.2+0.2\times\min\!\bigl(\lvert w_{i}-w_{j}\rvert/800,\;1\bigr)$: same family, different weight, where $w_{i},w_{j}\in\{100,200,\ldots,900\}$ are CSS font-weight values and the denominator 800 is the maximum possible weight difference (e.g., 400 vs. 500 yields $d=0.225$; 100 vs. 900 yields $d=0.4$).
  • $d=0.7$: different family, same typographic category (e.g., both sans-serif).
  • $d=1.0$: different typographic category (e.g., serif vs. sans-serif).
This distance function is symmetric and assigns zero distance only to identical variants and is fully determined by font metadata—it does not depend on the model’s learned representations, avoiding the circularity that would arise from using the model’s own embeddings to grade its errors.
@@
Table 2: Error severity analysis. Family-level accuracy collapses weight variants; relative severity $\Pi$ is a scale-invariant ratio of the model’s typographic-distance-weighted error to the random baseline.
@@
Top-1 accuracy (variant-level) 99.0%
Family-level accuracy 99.4%
SWER (model) 0.0054
SWER (random baseline) 0.7778
Relative severity $\Pi$ 0.0069
Figure 7 visualizes the font family similarity hierarchy recovered by UPGMA clustering on cosine distances between mean [CLS] embeddings. Note that this dendrogram uses model-derived distances (to show what the model has learned), whereas the SWER metric above uses the metadata-based typographic distance (to provide a model-independent evaluation). Serif, sans-serif, monospace, and display families occupy distinct branches in both, confirming that the model’s learned representations align with typographic structure.
6.3 Baseline Comparisons
To contextualize the efficiency–accuracy trade-off of our LoRA approach, we compare against standard fine-tuning strategies applied to the same DINOv2-Base backbone and dataset. Table 3 reports the results.
Table 3: Comparison of fine-tuning strategies on DINOv2-Base with 394 font classes. All methods are evaluated at 50 training epochs using the same training data, hyperparameters (where applicable), and evaluation protocol.
Method Trainable Params Top-1 Acc 95% CI $\Pi$
ResNet-50 (CNN) 25.6M 98.8% $\pm$0.17% 0.005
DINOv2-Base backbone:
Linear Probe 606K 36.6% $\pm$0.75% 0.479
LoRA (r=4r{=}4) 753K 97.9% $\pm$0.21% 0.013
LoRA (r=8r{=}8) 900K 98.7% $\pm$0.16% 0.007
LoRA (r=16r{=}16) 1.2M 98.7% $\pm$0.16% 0.007
Full Fine-Tuning 87.2M 93.4% $\pm$0.31% 0.028
[…]
We attribute full fine-tuning’s underperformance (93.4% at 50 epochs, 95.9% converged) to overfitting: with all 87.2M parameters trainable, the model has far more capacity than needed for $\sim$}575 training images per class, leading to degradation of the pretrained features that LoRA preserves by construction. Figure 9 provides direct evidence: the generalization gap (validation loss minus training loss) widens substantially for full fine-tuning over the course of training, while LoRA r=8r{=}8 maintains a narrow gap throughout.
Figure 8 shows the validation accuracy over training for each method. The LoRA configurations converge rapidly, reaching 95% accuracy within the first 5 epochs and plateauing near 99% by epoch 30. Full fine-tuning converges more slowly, requiring approximately 80 epochs to reach its final accuracy of 95.9%. The linear probe converges quickly but to a much lower ceiling, confirming that the frozen DINOv2 features alone are insufficient for fine-grained font discrimination. These convergence profiles demonstrate that LoRA’s parameter efficiency translates directly to faster training convergence, not just fewer trainable parameters.
[…]
Figure 8: Validation accuracy over training epochs for each fine-tuning strategy. LoRA variants converge to >>98% within 10 epochs, while full fine-tuning requires $\sim$80 epochs to reach 96%. The dashed line marks the 50-epoch evaluation point used in Table 3.
Figure 9: Left: training and validation loss for LoRA r=8r{=}8 and full fine-tuning. Right: generalization gap (validation loss minus smoothed training loss). Full fine-tuning exhibits a widening gap indicative of overfitting, while LoRA maintains a narrow gap throughout training. The briefly negative gap early in LoRA training is expected: dropout is active during training but disabled during evaluation, so validation loss can temporarily fall below training loss.
6.4 Generalization to Real-World Images
Our model is trained and evaluated entirely on synthetically rendered text. The most established real-world font recognition benchmark, AdobeVFR (Wang et al., 2015), contains 4,384 expert-labeled images spanning 617 font classes—however, these are drawn almost exclusively from commercial typefaces (Adobe Type Library, Linotype, Monotype), with zero overlap with our 32 Google Fonts families. Direct comparison is therefore not possible without retraining on a shared label space.
This highlights a gap in the font recognition literature: no public real-world benchmark exists for the open-source fonts that dominate modern web and mobile design. We are assembling a manually curated evaluation set of web typography screenshots for our target fonts, leveraging the fact that Google Fonts usage metadata identifies which websites use each family. Results on this real-world benchmark will be reported in a future revision. We note that our model is deployed in a production design tool and performs well on web screenshots in practice, though we refrain from reporting anecdotal production metrics as a substitute for controlled evaluation.
@@ 7 Deployment
The trained model is released on HuggingFace Hub as dchen0/font_classifier_v4.444https://huggingface.co/dchen0/font_classifier_v4 Before upload, the LoRA adapters are merged into the base weights (see Section 4.2), so the published checkpoint requires no additional libraries beyond HuggingFace Transformers. A custom handler.py is included for HuggingFace Inference Endpoints, accepting raw image bytes or base64-encoded strings and returning top-5 predictions with softmax confidence scores.
@@ 8 Discussion
Natural extensions include: (1) training on a broader set of fonts beyond the current 32 families, (2) incorporating real-world training images via domain adaptation, (3) jointly predicting font family and attributes (weight, width, style) in a multi-task framework, and (4) integrating text detection to enable end-to-end font recognition from uncropped images.
@@ 9 Conclusion
We have introduced GoogleFontsBench, a public benchmark for open-source font classification that fills a gap in existing evaluation resources dominated by commercial typefaces. On this benchmark, LoRA-adapted DINOv2 achieves 99.0% top-1 accuracy while training only 1% of parameters, with typographically mild errors ($\Pi=0.007$). The benchmark, baseline results from six methods, evaluation protocol, and full training pipeline are publicly available to support further research in font recognition and fine-grained visual classification. [^5] Source code: https://github.com/Create-Inc/font-model